a metal recycling plant has some scrap metal that is 44% copper.it also has some other scrap meaflthat is 29% copper the scrap metal is melted together and produces 1510g of metal that is 41% copper how many grams of the 29% copper the scrap metal were mixed with the 44% copper
Answer by mananth(16946)

Percent ---------------- quantity
Copper scrap I 44.00% ---------------- x gm
Copper scrap II 29.00% ------ 1510 - x gm
Mixture 41.00% ---------------- 1510
Total 1510 gm
44.00% x + 29.00% ( 1510 - x ) = 41.00% * 1510
44 x + 29 ( 1510 - x ) = 61910
44 x + 43790 - 29 x = 61910
44 x - 29 x = 61910 - -43790
15 x = 18120
/ 15
x = 1208 gm 44.00% Copper scrap I
302 gm 29.00% Copper scrap II